A first order reaction is 75% completed in 60min.find half-life period of the reaction.

Grade:Upto college level

2 Answers

Gaurav
9 years ago
Order of reaction = 1
Let the initial concentration be [A]0 = 100
[A]t = 100 -75 = 25
Rate law for first order rection
K = (2.303/t)log[A]0/[A]t
Substituting the values we get
K = 0.0231
And
t1/2 = 0.693/K = 30 minuts
Yash Chourasiya
3 years ago
Dear Student

Given
Order of reaction = 1
Let the initial concentration be [A]0 = 100
[A]t = 100 -75 = 25
Rate law for first order rection K = (2.303/t)log[A]0/[A]t
Substituting the values we get K = 0.0231
And t1/2 = 0.693/K = 30 minutes

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya